OpenStudy (anonymous):

5 sqrt 3 times 4 sqrt 15

OpenStudy (anonymous):

\[5 \sqrt3 \times 4 \sqrt15\]\[20 \sqrt45\]You multiply the numbers outside the root and the numbers inside the root separately. You can simplify this to...\[20\sqrt{5 \times 3 \times 3}\]\[60 \sqrt5\] Because the two 3s in the root can be taken out.